# Efficient and Accurate Optimal Transport with Mirror Descent and Conjugate Gradients

### v3

@@ B.1 Background: Line Search
Given a descent direction ${\bm{p}}^{(k)}\in\mathbb{R}^{n}$, i.e., a direction that satisfies $\langle{\bm{p}}^{(k)},\nabla f({\bm{x}}^{(k)})\rangle\leq 0$, line search algorithms aim to find an appropriate step size $\alpha$, where ${\bm{x}}^{(k+1)}\leftarrow{\bm{x}}^{(k)}+\alpha{\bm{p}}^{(k)}$. Perhaps the most well-known of desirable properties that a step size $\alpha$ should satisfy at any given optimization step are the Wolfe conditions (Wolfe, 1969, 1971). Given $\phi(\alpha)\coloneqq f({\bm{x}}^{(k)}+\alpha{\bm{p}}^{(k)})$:
@@
where $10<c_{1}<c_{2}<$ 1 and (28a) and (28b) are known as the sufficient decrease and curvature conditions respectively (Nocedal & Wright, 2006). It is well-known that given step sizes satisfying the Wolfe conditions and descent directions ${\bm{p}}^{(k)}$ that are not nearly orthogonal to the steepest descent directions $-\nabla f({\bm{x}}^{(k)})$, line search methods ensure convergence of the gradient norms to zero (Zoutendijk, 1966; Wolfe, 1969, 1971). Instead of satisfying (28), some algorithms or theoretical analyses consider exact line search, where ${\alpha^{*}\in\operatorname*{arg\,min}_{\alpha\in\mathbb{R}}\phi(\alpha)}$, which has a unique closed-form solution for quadratic objectives with a positive definite Hessian. However, a rule of thumb for general non-linear objectives is to not spend too much time finding $\alpha^{*}$ (Nocedal & Wright, 2006). Hager & Zhang (2006a) proposed approximate Wolfe conditions, derived by replacing the $\phi(\alpha)$ and $\phi(0)$ terms in (28a) with $q(\alpha)$ and $q(0)$, where $q$ is a quadratic interpolant of $\phi$ such that $q(0)=\phi(0)$, $q^{\prime}(0)=\phi^{\prime}(0)$ and $q^{\prime}(\alpha)=\phi^{\prime}(\alpha)$:
@@ B.2 PNCG Line Search
C Entropy-aware Stopping Criteria on the Dual Objective Gradient Norm
@@
D Variable vs. Fixed Smoothing of the Marginals in MDOT
As discussed in Sec. 4.2, MDOT smoothes the marginals ${\bm{r}},{\bm{c}}$ (by mixing in the uniform distribution) with a weighting factor that tracks the temperature. Since MDOT anneals the temperature, this means that the smoothing weight is higher in earlier iterations of MDOT. In particular, the mixture weight gradually decays from $H_{\min}({\bm{r}},{\bm{c}})\big{/}4\gamma_{\mathrm{i}}^{p}$ to $H_{\min}({\bm{r}},{\bm{c}})\big{/}4\gamma_{\mathrm{f}}^{p}$ given input parameter $p\geq 1$. Here, we study the effect of this design choice as it influences the convergence of two Bregman projection algorithms used in L7 of Alg. 1: Sinkhorn iteration and the newly proposed PNCG algorithm (Alg. 2). The approach is benchmarked against a baseline that fixes the smoothing weight at $H_{\min}({\bm{r}},{\bm{c}})\big{/}4\gamma_{\mathrm{f}}^{p}$ all throughout instead. For these experiments, we fix $p=1.5$ following our experimental setup in Sec. 5. Fig. 7 shows that while MDOT-Sinkhorn is largely unaffected by this design choice, MDOT-PNCG enjoys a notable speedup from variable smoothing. We thus conclude that the approach provides a performance benefit.
E Details of Baseline Algorithm Implementations
@@
Sinkhorn (Cuturi, 2013). A log-domain stabilized implementation was used. For different precision levels, $\gamma$ was varied from $2^{5}$ to $2^{14}$ for $L_{1}$ distance cost and to $2^{15}$ for $L_{2}^{2}$ distance cost. Stopping criteria were given by our formula in L4 of Alg. 1, and the results obtained by calling Alg. 1 with $\gamma_{\mathrm{i}}=\gamma_{\mathrm{f}}$, so that the algorithm terminates after a single Bregman projection via SK iteration.
@@
F Additional Benchmarking on DOTmark
Figs. 8-17 add further benchmarking on 10 more datasets from the DOTmark benchmark of Schrieber et al. (2017), which include various kinds of randomly generated images, classical test images and real data from microscopy. Each dataset contains 45 unique pairs of marginals $({\bm{r}},{\bm{c}})$ obtained from pixel values. The cost matrix is constructed from distances in 2D pixel locations; we evaluate on both $L_{1}$ and $L_{2}$ distance costs for $n=4096$ following our setup in Sec. 5.

### v4

@@ B.1 Background: Line Search
Given a descent direction ${\bm{p}}^{(k)}\in\mathbb{R}^{n}$, i.e., a direction that satisfies $\langle{\bm{p}}^{(k)},\nabla f({\bm{x}}^{(k)})\rangle\leq 0$, line search algorithms aim to find an appropriate step size $\alpha$, where ${\bm{x}}^{(k+1)}\leftarrow{\bm{x}}^{(k)}+\alpha{\bm{p}}^{(k)}$. Perhaps the most well-known of desirable properties that a step size $\alpha$ should satisfy at any given optimization step are the Wolfe conditions (Wolfe, 1969; 1971). Given $\phi(\alpha)\coloneqq f({\bm{x}}^{(k)}+\alpha{\bm{p}}^{(k)})$:
@@
where $10<c_{1}<c_{2}<$ 1 and (35a) and (35b) are known as the sufficient decrease and curvature conditions respectively (Nocedal & Wright, 2006). It is well-known that given step sizes satisfying the Wolfe conditions and descent directions ${\bm{p}}^{(k)}$ that are not nearly orthogonal to the steepest descent directions $-\nabla f({\bm{x}}^{(k)})$, line search methods ensure convergence of the gradient norms to zero (Zoutendijk, 1966; Wolfe, 1969; 1971). Instead of satisfying (35), some algorithms or theoretical analyses consider exact line search, where ${\alpha^{*}\in\operatorname*{arg\,min}_{\alpha\in\mathbb{R}}\phi(\alpha)}$, which has a unique closed-form solution for quadratic objectives with a positive definite Hessian. However, a rule of thumb for general non-linear objectives is to not spend too much time finding $\alpha^{*}$ (Nocedal & Wright, 2006). Hager & Zhang (2006a) proposed approximate Wolfe conditions, derived by replacing the $\phi(\alpha)$ and $\phi(0)$ terms in (35a) with $q(\alpha)$ and $q(0)$, where $q$ is a quadratic interpolant of $\phi$ such that $q(0)=\phi(0)$, $q^{\prime}(0)=\phi^{\prime}(0)$ and $q^{\prime}(\alpha)=\phi^{\prime}(\alpha)$:
@@ B.2 PNCG Line Search
Appendix C Entropy-aware Stopping Criteria on the Dual Objective Gradient Norm
@@
Appendix D Variable vs. Fixed Smoothing of the Marginals in MDOT
As discussed in Sec. 4.2, MDOT smoothes the marginals ${\bm{r}},{\bm{c}}$ (by mixing in the uniform distribution) with a weighting factor that tracks the temperature. Since MDOT anneals the temperature, this means that the smoothing weight is higher in earlier iterations of MDOT. In particular, the mixture weight gradually decays from $H_{\min}({\bm{r}},{\bm{c}})\big{/}4\gamma_{\mathrm{i}}^{p}$ to $H_{\min}({\bm{r}},{\bm{c}})\big{/}4\gamma_{\mathrm{f}}^{p}$ given input parameter $p\geq 1$. Here, we study the effect of this design choice as it influences the convergence of two KL projection algorithms used in L6 of Alg. 1: Sinkhorn iteration and the newly proposed PNCG algorithm (Alg. 2). The approach is benchmarked against a baseline that fixes the smoothing weight at $H_{\min}({\bm{r}},{\bm{c}})\big{/}4\gamma_{\mathrm{f}}^{p}$ all throughout instead. For these experiments, we fix $p=1.5$ following our experimental setup in Sec. 5. Fig. 8 shows that while MDOT-Sinkhorn is largely unaffected by this design choice, MDOT-PNCG enjoys a notable speedup from variable smoothing. We thus conclude that the approach provides a performance benefit.
Appendix E Comparison with CPU-based Solver for Higher $n$
In Figure 5, we displayed via a vertical line the time taken on a CPU by the network simplex solver from the Python Optimal Transport package of Flamary et al. (2021). Clearly, a faster CPU would benefit the network simplex and a faster GPU would benefit the MDOT family of algorithms, which makes them difficult to compare in a fair setup. However, given the known $\widetilde{O}(n^{3})$ complexity of the network simplex and the $\widetilde{O}(n^{2})-\widetilde{O}(n^{2.5})$ empirical dependence of MDOT-PNCG seen in Fig. 6, we suspect that the precision-speed trade-off posed by MDOT-PNCG improves with increasing $n$. In Table 1, we provide a comparison of the two algorithms, with several values of $\gamma_{\mathrm{f}}$ used for MDOT-PNCG and $n$ increased from $4,096$ to $16,384$. We observe the same trend across both $L_{1}$ and $L_{2}$ cost functions on the MNIST dataset; for a comparable level of relative error achieved by MDOT–PNCG (at a fixed $\gamma_{\mathrm{f}}$), the speedup over the network simplex solver is better for $n=16,384$. We expect the trend to continue with higher $n$, and note that MDOT stands to benefit from ongoing rapid developments in GPU hardware innovation, as well as faster projection algorithms than PNCG; see also our concurrent work in this direction showing substantial speedups (Kemertas et al., 2025).
Table 1: Comparison of the exact Network Simplex solver with MDOT–PNCG method on MNIST transport problems under $L_{1}$ and $L_{2}$ costs. Relative error is computed as $100*\langle P-P^{*},C\rangle/\langle P^{*},C\rangle$.
$n=4,096$ $n=16,384$
Cost Fn. Algorithm $\gamma_{\mathrm{f}}$ RelErr % Time (s) Speedup RelErr % Time (s) Speedup
${}~{}~{}L_{1}$ Net. Simplex 0.0000 5.45 1.00 0 236.46 1.00
MDOT–PNCG $2^{6}$ 16.556 0.18 30.55 22.580 1.72 137.24
$2^{9}$ 0.167 2.28 2.39 0.585 26.94 8.78
$2^{12}$ 0.002 11.85 0.46 0.002 318.58 0.74
${}~{}~{}L_{2}^{2}$ Net. Simplex 0.000 11.74 1.00 0 728.40 1.00
MDOT–PNCG $2^{9}$ 26.877 0.52 22.66 23.827 5.53 131.83
$2^{12}$ 3.166 3.85 3.04 3.372 38.12 19.11
$2^{15}$ 0.044 39.38 0.30 0.303 294.12 2.48
Appendix F Details of Baseline Algorithm Implementations
@@
Sinkhorn (Cuturi, 2013). A log-domain stabilized implementation was used. For different precision levels, $\gamma$ was varied from $2^{5}$ to $2^{14}$ for $L_{1}$ distance cost and to $2^{15}$ for $L_{2}^{2}$ distance cost. Stopping criteria were given by our formula in L4 of Alg. 1, and the results obtained by calling Alg. 1 with $\gamma_{\mathrm{i}}=\gamma_{\mathrm{f}}$, so that the algorithm terminates after a single KL projection via SK iteration.
@@
Figure 9: Optimality gap $\langle P-P^{*},C\rangle$ of rounded (feasible) plans vs. number of $O(n^{2})$ operations for IPOT (Xie et al., 2020) and MDOT (with Sinkhorn and PNCG for KL projections, using $p=1.5,q=2^{1/3}$ as in Section 5). Both methods are run up to a final $\gamma_{\mathrm{f}}=2^{15}$ on the upsampled MNIST dataset ($n=4096$) with $L_{1}$ (left) and $L_{2}^{2}$ (right) cost functions. MDOT reaches the target $\gamma_{\mathrm{f}}$ satisfying its stopping criterion $\left\lVert\nabla g(\gamma_{\mathrm{f}})\right\rVert_{1}=O(\gamma^{-p})$ in fewer operations than IPOT takes until termination, so we continue the KL projection at $\gamma_{\mathrm{f}}$ until the same number of operations as IPOT is reached, without increasing $\gamma$ further (dashed lines).
Appendix G Similarities and Differences with IPOT
In this section, we discuss the similarities and differences between MDOT and the Inexact Proximal point method for exact Optimal Transport (IPOT) algorithm of Xie et al. (2020) in more detail. In IPOT, authors propose to update (in our notation) the $\gamma$ parameter by fixed increments and typically choose $\Delta_{\gamma}^{(t)}=1$ for all $t\geq 0$, while we took $\Delta_{\gamma}=(q-1)\gamma$ for a hyperparameter ${q>1}$. Following each temperature update, they run a fixed number of $L$ row+column scaling (Sinkhorn) updates and recommend ${L=1}$, whereas MDOT requires that the KL projection is continued until the dual objective gradient norm reaches below a threshold at each $\gamma$. A small number of $L$ Sinkhorn updates clearly does not amount to an “exact” KL projection onto the feasible set ${\mathcal{U}}({\bm{r}},{\bm{c}})$, but Xie et al. (2020) show that under some conditions there exists some finite $L$ that guarantees linear convergence of $\langle P({\bm{u}},{\bm{v}};\gamma),C\rangle$ to the $\langle P^{*},C\rangle$. On the other hand, (i) MDOT naturally adapts the number of optimization updates to the difficulty of the problem and the strictness of the stopping criterion at $\gamma$, and (ii) it stands to benefit from potentially faster convex optimization algorithms besides Sinkhorn iteration, e.g., our PNCG approach proposed in Sec. 4 or any other approach that may be developed in the future. A similarity between the approaches is that IPOT also includes an implicit warm-start of the dual variable, which can be considered a special case of our warm-start proposed in Section 4.2.1, where $\Delta_{\gamma}^{(t)}=1$ for all $t$.
In Figure 9, we compare IPOT with MDOT ($p{=}1.5,q{=}2^{1/3},\gamma_{\mathrm{i}}{=}2$) on the upsampled MNIST dataset. For a fair comparison, we followed the conventions of Xie et al. (2020) and implemented all algorithms by computing row/column sums of $P({\bm{u}},{\bm{v}};\gamma)$ explicitly rather than via LogSumExp reductions as in Algorithms 2 and 4.444In this implementation, we evaluate the Sinkhorn direction (15) for PNCG by adding a small constant ${\approx}10^{-30}$ to each entry of ${\bm{r}}(P)$ and ${\bm{c}}(P)$ for numerical stability. Each matrix-vector product, row/column sum, row/column scaling of a matrix and entry-wise operations on matrices is counted as one operation of cost $O(n^{2})$. Since MDOT satisfies its stopping criterion (reaches $\gamma_{\mathrm{f}}$ and satisfies $\left\lVert\nabla g(\gamma_{\mathrm{f}})\right\rVert_{1}\leq H_{\min}({\bm{r}},{\bm{c}})/\gamma^{-p}$) up to 10×10$\times$10 × more quickly, we continue minimizing $g({\bm{u}},{\bm{v}};\gamma,{\bm{r}},{\bm{c}})$ using the respective KL projection algorithm (Sinkhorn or PNCG) until the same number of $O(n^{2})$ operations as IPOT are executed (namely, $5\times\gamma_{\mathrm{f}}$).
Overall, the methods behave similarly for the $L_{1}$ cost, while MDOT is up to 10×10$\times$10 × faster for the $L_{2}^{2}$ cost.
These results also reveal an interesting contrast between the $L_{1}$ and $L_{2}^{2}$ cost functions. For the $L_{1}$ cost the optimality gap rapidly approaches 0 as the KL projection becomes more precise (see dashed lines), even though $\gamma_{\mathrm{f}}$ is fixed. This suggests that the entropic gap, namely $\langle P^{*}(\gamma)-P^{*},C\rangle$, is small at $\gamma_{\mathrm{f}}$ for these MNIST problems, and the optimality gap is dominated by the inexactness of the projection onto ${\mathcal{U}}({\bm{r}},{\bm{c}})$. We believe that the fast rate $O(\exp\{-\gamma K\})$ of Weed (2018) discussed in Section 4.2.2 is active here for the entropic gap, although this requires further study. On the other hand for the $L_{2}^{2}$ cost, continuing to iterate on the projection error to improve the approximation of $P^{*}(\gamma_{\mathrm{f}})$ does not reduce the optimality gap any further, which implies the entropic gap $\langle P^{*}(\gamma_{\mathrm{f}})-P^{*},C\rangle$ is still dominant at $\gamma_{\mathrm{f}}$ in these problems.
Appendix H Additional Benchmarking on DOTmark
Figs. 10-19 add further benchmarking on 10 more datasets from the DOTmark benchmark of Schrieber et al. (2017), which include various kinds of randomly generated images, classical test images and real data from microscopy. Each dataset contains 45 unique pairs of marginals $({\bm{r}},{\bm{c}})$ obtained from pixel values. The cost matrix is constructed from distances in 2D pixel locations; we evaluate on both $L_{1}$ and $L_{2}$ distance costs for $n=4096$ following our setup in Sec. 5.